The longer leg of a right triangle measures 5 inches less than the hypotenuse. The shorter leg measures 25 inches. Find the length of the long leg.

Answer by scott8148(6628)

let x="longer leg", so hypotenuse is x+5

by Pythagoras __ x^2+25^2=(x+5)2 __ x^2+625=x^2+10x+25

subtracting x^2+25 __ 600=10x __ x=60